OpenStudy (anonymous):

What is the sum of the arithmetic sequence 137, 125, 113 …, if there are 38 terms?

myininaya (myininaya):

Do you see that the numbers are increasing by 8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i would do it the old school way and keep subtracting 12 until you get 38 terms, then add lol

myininaya (myininaya):

\[\sum_{i=0}^{37}(137+8i)=137+8(0)+\sum_{i=1}^{37}(137+8i)\] Now use \[\sum_{i=1}^{n}i=\frac{n(n+1)}{2} \text{ and } \sum_{i=1}^{n}c=cn\]

OpenStudy (anonymous):

Im still not getting the correct answer

OpenStudy (zarkon):

whay are people using d=8

OpenStudy (zarkon):

*why

OpenStudy (anonymous):

i dont know im gonna guess

OpenStudy (anonymous):

Whenever you need to find the sum of an arithmetic series use this formula: \[ \frac n 2 \left( 2a+ (n-1) \right)\times d) \tag{1}\] a= first term n= number of terms d= common difference Here, \(n= 38, a = 137 \) and \( d = -12\). Substituting these in \((1)\), and assuming my algebra is right you answer should be \(-3230 \)

OpenStudy (anonymous):

I know but the answer does not show up in the multiple choose

OpenStudy (anonymous):

@NatalieLove: Revert if you need help in proving that formula.

OpenStudy (anonymous):

We are doing the same steps the test its incorrect. It does not show up as a choose Thank You

OpenStudy (anonymous):

Glad to help :)

myininaya (myininaya):

omg lol I said increasing by 8 and then continue to think the thingy was increasing by 8 sometimes i wonder about myself lol